OpenStudy (anonymous):

Simplify the following expression: 3(3z - 4 ) - (20 z - 5)

mathslover (mathslover):

first of all expand the brackets: \[\large{3(3z-4)-(20z-5)=3(3z)-3(4)-20z-(-5)}\]

mathslover (mathslover):

now we have : \[\large{9z-12-20z+5}\] right?

mathslover (mathslover):

@julialeigh please reply

OpenStudy (anonymous):

yeah how exactly did u get the 9z-12-20z+5

mathslover (mathslover):

what is 3*3z ?

OpenStudy (anonymous):

9

mathslover (mathslover):

where did z go?

OpenStudy (anonymous):

after the 9 now the 12

mathslover (mathslover):

what is 3*x ?

OpenStudy (anonymous):

ok got that it would be 3z times that (4) right

OpenStudy (anonymous):

then -20z +5.. so it would be 9z-12-20z+5

OpenStudy (anonymous):

plz reply

mathslover (mathslover):

right

mathslover (mathslover):

sorry for late reply

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

how do i simplify that

mathslover (mathslover):

9z - 20z = -11 z -12 + 5= -7 so we have : \[\large{-11z-7}\]

OpenStudy (anonymous):

thank you